OpenStudy (anonymous):

Add. 36 + 40 + (–2) A. –78 B. –74 C. 74 D. 78

OpenStudy (anonymous):

@elcna

OpenStudy (texaschic101):

36 + 40 + (-2) --- a positive times a negative = a negative 36 + 40 - 2 = yo can do the math now

OpenStudy (anonymous):

C

OpenStudy (anonymous):

36 + 40 = 76 76 - 2 = 74